OpenStudy (anonymous):

Penelope is making a quilt that has 4 sections, each of which is 2 5/8 feet in length. What is the length of the entire quilt?

OpenStudy (goldphenoix):

So 1 section is 2 5/8 feet long?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you plz solve it for me

OpenStudy (goldphenoix):

Sure. So if: 1 section is 2 5/8 feet long. Then 4 section is... What do you think you have to do?

OpenStudy (anonymous):

do 2 5/8 x 4?

OpenStudy (goldphenoix):

Correct!

OpenStudy (goldphenoix):

Do you know how to multiply fraction and whole number?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Hello you there?

OpenStudy (goldphenoix):

If not, then convert 2 5/8 into an improper fraction.  ^ That is how you convert it into an improper fraction. You times the denonimator by the whole number and add it by the numerator. So 2 x 8 + 5 = 21. The denonimator stays the same. So 2 5/8 in improper form would be 21/8. How do you times 21/8 x 4? \[\frac{ 21 }{ 8 } \times 4\] It's easy, you can convert 4 into an improper fraction by putting 1 as a denonimator. So 4/1 is the improper fraction of 4! Now it should be: \[\frac{ 21 }{ 8 } \times \frac{ 4 }{ 1 }\] Time it regularlly. 21 x 4 is equal to 84. 8 x 1= 8. So the fraction would be: \[\large \frac{ 84 }{ 8 }\] You want to convert it back into mixed number. To do this, you must divide the numerator by the denonimator. So 84 divided by 8 will give you 10.5 So the length of the entire quilt is 10.5 feet! I hope this help!